OpenStudy (anonymous):

2^7x=83 using logarithmsss

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow log_{2}83 = 7x }\) Do you know about logarithms?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x^n = y \Longleftrightarrow log_{x}y = n }\)

OpenStudy (lgbasallote):

\[\LARGE x = \frac{\log_2 83}{7} = \frac{1}{7} \log_2 83 = \log_2 83^7\] just to make it easier

OpenStudy (anonymous):

@lgbasallote \[\huge{\log_283^{\frac{1}{7}}}\]

OpenStudy (lgbasallote):

oh yes...thanks for the correction

OpenStudy (lgbasallote):

\[\LARGE \log_2 83^{\frac{1}{7}} = \log_2 \sqrt[7]{83}\]

OpenStudy (anonymous):

I have to solve for x...im not sure if this is right but i have [x=\log83-7\log_2/\log_2\] what do i do nextt??

OpenStudy (lgbasallote):

\[x=\log83-7\log_2/\log_2\]

OpenStudy (lgbasallote):

how do you read that?

OpenStudy (anonymous):

yes, just like that

OpenStudy (lgbasallote):

i don't get the \(7\log_2/\log_2\) part

OpenStudy (anonymous):

log_2 is on the bottom the rest is the numerator

OpenStudy (lgbasallote):

well it's impossible to just have \(7\log_2\) you need a power

OpenStudy (lgbasallote):

let me just tell you that \[\LARGE \frac{\log_2 83}{7} \ne \log_2 (\frac{83}{7})\]

OpenStudy (lgbasallote):

\[\LARGE \frac{\log_2 83}{7} = \log_2 83^{\frac{1}{7}}\] \[\mathbf{BUT}\] \[\LARGE \log_2 (\frac{83}{7}) = \log_2 83 - \log_2 7\] therefore they are NOT equal

OpenStudy (anonymous):

nooo lgbasallote, thats not the equation but i dont quite know how to write it on this srinidhijha! That looks right but why did you write 7xLog_2, its not just 7xLog ?

OpenStudy (lgbasallote):

i dont know how you got \(7\log_2\) :/

OpenStudy (anonymous):

@Lovely95 bcz i have taken log on boyh sides with base 10 . If u dont write a base ex-2,e then it means base of log is 10 so 7xlog2 = (7x ) {log base 10 (2)} whichz not equal to (7x) log base 10

OpenStudy (anonymous):

@srinidhijha But .228 doesnt check out :/

OpenStudy (anonymous):

mistaken! this is the correct answer you can do it by taking base 2 . its simpler and answer is CORRECT for sure

OpenStudy (anonymous):

Yes! its checks THANK YOUU